OpenStudy (anonymous):

Solve for x: -3x = 11

OpenStudy (anonymous):

i believe the answer is x= -3/11 but i wanted to check and make sure

jimthompson5910 (jim_thompson5910):

let's check that answer

jimthompson5910 (jim_thompson5910):

\[\large -3x = 11\] \[\large -3\left(-\frac{3}{11}\right) = 11 ... \ \text{Plug in x = } -\frac{3}{11}\] \[\large \left(-\frac{3}{1}\right)\left(-\frac{3}{11}\right) = 11\] \[\large \frac{-3*(-3)}{1*11} = 11\] \[\large \frac{9}{11} = 11\] is that last equation true?

OpenStudy (anonymous):

i dont think so

jimthompson5910 (jim_thompson5910):

the answer is no because you want 11 = 11

jimthompson5910 (jim_thompson5910):

so you want the left side to turn into 11

OpenStudy (anonymous):

what do you get if you divide both sides of equation by -3 ???

OpenStudy (anonymous):

x=-11/3 is the answer if im not mistaken

jimthompson5910 (jim_thompson5910):

go ahead and check it to be sure